OpenStudy (mathmath333):

What annual payment will discharge a debt of Rs.808 due in 2 years at 2% per annum ? a. 200 b. 300 c. 400 d. 350

OpenStudy (anonymous):

so u need to pay 2% of 808 for first year and pay 2%of (2% of 808) for second year

hartnn (hartnn):

compound interest, right? \(\Large A =P (1+(1+r)^n)\)

hartnn (hartnn):

A=808 r = 0.02 n=2 find P

OpenStudy (anonymous):

hehe i dont memorize i just analyze them :)

OpenStudy (mathmath333):

why are'nt you using simple interest

hartnn (hartnn):

like A = P(1+rn) ?

OpenStudy (mathmath333):

i=pr%t

hartnn (hartnn):

ok, what answer does it give you ? I use compound Interest, because I assume, the rate of interest applies on new amount after every year...

OpenStudy (mathmath333):

book is giving \(400\)

hartnn (hartnn):

with compound Interest we are getting 396

ganeshie8 (ganeshie8):

suppose M is the required yearly payment

ganeshie8 (ganeshie8):

808 in 2 years is same as 808/(1+0.02)^2 now so you want this amount to vanish to 0 by making a payment of M each year

ganeshie8 (ganeshie8):

balance after 1 year : \[\dfrac{808}{(1+0.02)^2}(1+0.02) - M\]

ganeshie8 (ganeshie8):

balance after two years : \[\left[\dfrac{808}{(1+0.02)^2}(1+0.02) - M\right](1+0.02) -M\]

ganeshie8 (ganeshie8):

you want this balance to equal 0 as you're closing the loan set it equal to 0 and solve M

OpenStudy (mathmath333):

thnx both btw which concept are u using for this can u give a link